OpenStudy (anonymous):

someone please help me to solve this :( i want to learn!and i suck at geometry! Circle R has equation (x+2)^2+y^2=4. What is the center and radius of this circle?

OpenStudy (anonymous):

The center is (-2, 0) The radius is 2, I believe

OpenStudy (anonymous):

could you explain to me how you did that?

OpenStudy (anonymous):

So the basic equation of a circle is: (x-h)^2 + (y-k) =r^2 (h,k) is the center of the circle and r is the radius since is (x+2)^2, h must equal -2. You can figure this out by take the square root of this which makes it x+2=0 and just solve for x, which is h. You can do the same thing as for k, since there nothing there then it must be 0. Since it is 4 at the end with out it being squared then you must square root it. Does that make any sense

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

oh ok i understand how you got the center

OpenStudy (anonymous):

Okay, that's good

OpenStudy (anonymous):

but what would y equal?

OpenStudy (anonymous):

x and y are the places where you can plug in an order pair and see if that pair lies on the circle

OpenStudy (anonymous):

Ok thank you (: